# Detection of Intelligent Tampering in Wireless Electrocardiogram Signals Using Hybrid Machine Learning

### v1

@@ Abstract
With the proliferation of wireless electrocardiogram (ECG) systems for health monitoring and authentication, protecting signal integrity against tampering is becoming increasingly important. This paper analyzes the performance of CNN, ResNet, and hybrid (Transformer+CNN) models for tamper detection. It also evaluates the performance of a Siamese network for ECG-based identity verification. Six tampering strategies, including structured segment substitutions and random insertions, are emulated to mimic real-world attacks. The one-dimensional ECG signals are transformed into a two-dimensional representation in the time-frequency domain using the continuous wavelet transform (CWT). The models are trained and evaluated using ECG data from 54 subjects recorded in four sessions (2019–2025) outside of clinical settings while the subjects performed seven different daily activities. Experimental results show that in highly fragmented manipulation scenarios, CNN, FeatCNN-TranCNN, FeatCNN-Tran and ResNet models achieved an accuracy exceeding 99.5%. Similarly, for subtle manipulations (e.g., 50%–50%) and (75%–25%) substitutions, our FeatCNN-TranCNN model demonstrated consistently reliable performance, achieving an average accuracy of 98%. For identity verification, the pure Transformer-Siamese network achieved an average accuracy of 98.30%. In contrast, the hybrid CNN+Transformer-Siamese model delivered perfect verification performance with 100% accuracy.
@@ I Introduction
Electrocardiogram (ECG) signals have played an important role in the diagnosis and monitoring of cardiovascular diseases [1, 2]. As interest in preventive medicine grows, wireless ECG signals are being used for continuous health monitoring and biometric authentication [3, 4, 5]. However, wireless ECG data is vulnerable to different kinds of attacks [6, 7]. An ECG signal can be manipulated by replacing some part of ECG segment, or it can be manipulated by replacing the whole part of the ECG segment with another person’s ECG segment. Also, the manipulation can be conducted by using artificially generated ECG signals, and all these manipulations can affect the integrity of biometric systems [8, 9]. Several studies have been conducted to detect tampering or anomalies in ECG signals, ranging from traditional methods to machine learning-based systems [10]. Traditional methods identify manipulations, but they face scalability issues and struggle to identify complex attack strategies [11, 12, 13]. Conventional machine learning approaches, such as Support Vector Machines (SVMs) and Random Forests, have better performance compared to traditional methods [14, 15, 16]. However, these methods rely on manually created features, and because of this, they are less adaptive to unknown and more complicated attacks. Deep learning models, such as feed-forward networks and convolutional neural networks (CNNs), have been used to identify ECG signal manipulation, but they are not evaluated with sophisticated manipulation techniques [17, 18, 19, 20]. Likewise, Siamese Neural Networks [21, 22, 23] are employed for person-specific verification, relying on individual ECG features to improve robustness between subject variability.
In this paper, a comprehensive deep learning framework used for tampering detection as well as biometric identification is presented. The performance of different deep learning models, such as CNNs, ResNet, Transformer models, and hybrid models (Transformer+CNNs) are investigated under several tampering scenarios. In addition to this, the performance of a Siamese neural network in biometric verification is evaluated. The experiments are conducted on a multi-session ECG dataset of 54 subjects, and the results show that a ResNet model and a hybrid CNN-Transformer model achieved high accuracy in tamper detection and a hybrid CNN-Transformer model achieved high accuracy in identity verification. The accuracy of CNN and ResNet models is high in detecting highly localized tampering; however, the training results show that these models highly depend on specific training parameters. The hybrid Transformer-CNN based models achieved high accuracy as well as a high degree of stability and consistent training results. The study conducted highlights that the hybrid Transformer-CNN model is more effective at securing wireless ECG signals, thus improving the reliability of remote medical monitoring and biometric authentication. The contributions of this paper are summarized as follows:
@@
  • Data: Our approach is real-world data-driven. A large volume of ECG data was collected from 54 subjects carrying out 7 different activities. The data were collected in four batches between 2019 and 2025 and exhibit a high variance, both across subjects and across activities.
  • Model for recognizing tampering: Multiple machine learning frameworks such as CNN, ResNet, and hybrid models were proposed for tampering detection, and Siamese Neural Networks are proposed for robust identity verification. All these models were evaluated against different tampering strategies, and the performance of each model is compared with the others.
  • Experiments: Extensive and independent experiments were conducted, and the results show that the proposed models are effective in identifying complex ECG tampering. Specifically, hybrid models achieved high performance in tamper detection against various tampering attacks on wireless ECG data. Also, Siamese Networks achieved accuracy of more than 99% in person verification.
@@ II Related Work
The widespread use of wireless ECG data for healthcare monitoring and person identity verification has led to concern about the wireless ECG data’s authenticity. Various ECG data tampering detection mechanisms are proposed to improve the reliability of wireless ECG data, and the proposed approaches are discussed in this section.
@@
The proposed approaches demonstrate reliable detection of complex signal manipulations. However, most of these approaches account for external manipulations. In contrast, the present work does not rule out internal manipulation, i.e., the deliberate manipulation of ECG data by users to achieve various goals, such as concealing heart conditions and making insurance claims. In contrast to the data used in previous case studies, which were collected in controlled environments, the present work used data collected outside of clinical settings during various activities by subjects. Such data are difficult to evaluate and contain significant motion artifacts. The robust machine learning models we propose ensure that these challenges are successfully overcome.
@@ III Data Acquisition
2019 8 8/0 $30\pm 6$ years
2024 16 5/11 $27\pm 13$ years
2024 10 5/5 $22\pm 1.9$ years
2025 20 17/3 $24.2\pm 2.26$ years
@@ IV-E Tampering ECG Signal
Figure 5: Tampered ECG segment with linear blending.
@@ V-A3 Transformer Model
V-B Model Architecture for Person Verification
In this paper, a Siamese network architecture is used for person verification. The Siamese approach works by comparing two input sequences and drawing inferences related to their similarity. Each branch of the Siamese network operates on a single ECG input separately and identically, with every parameter being shared between the two branches so that embedding spaces are consistent. The network maps each input to a compressed latent form, to which a similarity score is calculated based on the distance between the two embeddings. Two encoders are studied under the Siamese framework, where each of them worked upon a different form of input data.
Figure 16: Siamese Network Model Deep Transformer Encoder
V-B1 Siamese Network with Deep Transformer Encoder
In the first variant, an ECG segment of shape 2048 $\times$ 96 is used as input, which corresponds to 4 s of the cardiac activity mapped into the time-frequency domain. The branch architecture adopts the deep Transformer encoder structure described earlier in Fig. 14, with the only change being the decrease of the number of attention heads from 8 to 4 in order to fit the scale of the verification task better, to a balance between expressive capacity and computational efficiency. The rest of the components, such as the multi-head self-attention mechanism, deep feed-forward network (FFN), residual connections, and positional encoding, are all in line with the original Transformer model.
V-B2 Siamese Network with CNN Feature Extractor and Transformer Encoder
In the second version, all branches process raw one-dimensional ECG inputs of the form 2048$\times$1. Hierarchical capture of localized morphological patterns is first obtained by a convolutional feature extractor; the extracted features are then fed to a Transformer encoder in a similar structure to the Transformer encoder discussed in Fig. 14. Empirical results verified that such a combination improves the person verification performance over a pure Transformer, indicating that multi-level feature aggregation, i.e., local via CNNs and global via self-attention, offers a more discriminative embedding space for ECG-based identity verification.
Figure 17: Siamese Network with CNN Feature Extractor and Transformer Encoder
V-B3 Training and Evaluation for Tampering Detection Models
@@
V-B4 Training and Evaluation for Person Verification Model
The models for person verification were trained on pairs of ECG segments, with two balanced categories: positive pairs segment, segments formed from the same individual but different activities, and negative pairs segment, segments formed from different individuals performing the same activity. This pairing strategy enabled the model to learn identity features independent of activity variations. The input format used by the two model are different, the Siamese-Transformer model received CWT segments of size 2048$\times$ 96, while the Siamese-CNN-Transformer model used raw 1D ECG signals of size 2048$\times$1.
V-C Computational Efficiency of Models
@@
TABLE II: Model Computational Effort and Input Size
Model Input Size FLOPs (Millions)
@@
VI Evaluation
For tampering identification, seven models are proposed, and the performance of each deep learning model is evaluated across six tampering scenarios. For person identity verification, two Siamese network models are evaluated against different attacks. Stratified sampling is used to split the data collected into 80% training, 10% validation, and 10% testing sets. Each experiment is repeated 25 times, and the reported results represent the average accuracy over these runs.
VI-A Tampering Detection Results
This section presents the performance of the seven deep learning models for detecting ECG tampering. As shown in Table, III, all models performed well in the classification of heavily tampered signals, especially in the scenarios with sporadic 20% or 50% tampering, where detection accuracies varied from 91% to 98% respectively. The performance differences between the individual models became more apparent in more subtle cases such as 50:50 and 75:25 segment swaps. The models combining convolutional feature extraction with transformer encoding achieved some of the best results. Interestingly, ResNet was consistently competitive in all scenarios, though the transformer-based models exhibited the smallest performance variance across the 25 independent experiments. The CNN-based models exhibited significant performance fluctuations.
@@ VI-A Tampering Detection Results
To further gain insights about the performance of each model, the performance matrices were separately analyzed for each model and summarized in Tables  IV through table  IX. Table  IV demonstrates that the 50-50 tampering was challenging for most models, while FeatCNN-TranCNN model obtained an accuracy of around 98.46%. Table  V demonstrate that the 75–25 tampering scenario presented the subtle and smoothly blended changes, and the FeatCNN-TransCNN and FeatCNN-Trans models obtained the highest accuracies of 98.7% and 98.5%, respectively. The 50–25–25 tampering scenario involved a combination of medium-sized segment replacements, and several models coped well with this tampering, as can be seen in table  VI. For 50% Alternating tampering shown in table  VII, the FeatCNN-Trans, FeatCNN-TransCNN and Resnet models had perfect detection performance, attaining 100% accuracy, precision, recall, and F1-score on all trials.
TABLE IV: Detailed Metrics (%) for 50-50 Tampering
@@
Tran-DeepFFN 85.48 84.29 86.69 85.47
Tran-CNNFFN 86.48 85.45 87.53 86.48
FeatCNN-Tran 98.07 98.02 98.13 98.07
FeatCNN-TranCNN 98.46 98.78 98.13 98.45
CWT-FeatCNN-Tran 94.88 94.35 95.41 94.88
CNN 97.90 97.73 98.05 97.90
ResNet 98.08 97.15 99.03 98.08
TABLE V: Detailed Metrics (%) for 75-25 Tampering
Model Accuracy Precision Recall F1-Score
Tran-DeepFFN 85.62 84.12 87.13 85.60
Tran-CNNFFN 85.80 84.74 86.87 85.79
FeatCNN-Tran 98.54 98.17 98.88 98.54
FeatCNN-TranCNN 98.68 98.46 98.90 98.68
CWT-FeatCNN-Tran 95.48 94.91 96.05 95.48
CNN 98.30 97.62 98.95 98.28
ResNet 98.78 98.50 99.05 98.78
TABLE VI: Detailed Metrics (%) for 50-25-25 Tampering
Model Accuracy Precision Recall F1-Score
Tran-DeepFFN 87.14 85.88 88.39 87.12
Tran-CNNFFN 87.30 86.59 87.98 87.28
FeatCNN-Tran 99.81 99.78 99.83 99.81
FeatCNN-TranCNN 99.78 99.81 99.78 99.78
CWT-FeatCNN-Tran 95.50 94.91 96.05 95.48
CNN 99.47 99.20 99.60 99.40
ResNet 98.36 97.77 98.95 98.36
TABLE VII: Detailed Metrics (%) for 50% Alternating Tampering
Model Accuracy Precision Recall F1-Score
Tran-DeepFFN 85.22 84.59 85.82 85.20
Tran-CNNFFN 92.10 91.49 92.67 92.08
FeatCNN-Tran 100.00 100.00 100.00 100.00
FeatCNN-TranCNN 100.00 100.00 100.00 100.00
CWT-FeatCNN-Tran 98.54 98.48 98.56 98.52
CNN 99.59 99.36 99.84 99.60
ResNet 100.00 100.00 100.00 100.00
TABLE VIII: Detailed Metrics (%) for Sporadic 20% Tampering
Model Accuracy Precision Recall F1-Score
CNN 99.80 99.84 99.76 99.80
ResNet 99.90 99.76 100.00 99.88
Tran-DeepFFN 91.33 90.70 91.87 91.28
Tran-CNNFFN 91.62 90.37 92.86 91.60
FeatCNN-Tran 99.89 100.00 99.78 99.89
FeatCNN-TranCNN 99.83 99.89 99.78 99.83
CWT-FeatCNN-Tran 99.36 99.12 99.44 99.28
TABLE IX: Detailed Metrics (%) for Sporadic 50% Tampering
Model Accuracy Precision Recall F1-Score
CNN 99.90 99.76 100.00 99.88
ResNet 100.00 100.00 100.00 100.00
Tran-DeepFFN 98.41 97.46 99.35 98.40
Tran-CNNFFN 98.91 98.33 99.44 98.88
FeatCNN-Tran 99.92 100.00 99.83 99.92
FeatCNN-TranCNN 99.94 100.00 99.89 99.94
CWT-FeatCNN-Tran 99.81 99.68 99.92 99.80
VI-B Person Specific Verification Results
Two types of Siamese models were developed for person identification, and the contrastive loss function is used to train the models. For variant 1, the Siamese Transformer ECG segment of shape 2048 $\times$ 96 is used, and for variant 2, the Siamese FeatureCNN-Transformer, one-dimensional ECG segments of the form 2048 $\times$ 1 is used. Pairing Strategy: Pairing strategy is conducted to make sure that the model learned person-specific characteristics rather than activity-specific patterns. As shown in table X ECG segments from the same individuals who perform different activities are used to form the positive pairs, and ECG segments from different individuals while performing the same activity are used to form the negative pairs.
TABLE X: Examples of Positive and Negative Pair
Pair Type S 1 (Activity) S2 (Activity) Label
Positive Pair S25 (Walking) S25 (Sitting) 1
Negative Pair S25 (Walking) S36 (Walking) 0
Positive Pair S40 (Cycling) S40 (Running) 1
Negative Pair S28 (Sitting) S47 (Sitting) 0
The model is evaluated on an unseen test set consisting of similarly constructed positive and negative pairs and summarized in table  XI.
TABLE XI: Detailed Metrics for Person Verification using Siamese Networks with Different Encoder Architectures
Encoder Used Accuracy (%) Precision (%) Recall (%) F1-Score (%)
Tran-DeepFFN 98.30 98.99 97.60 98.29
FeatCNN-Tran 100.00 100.00 100.00 100.00
VI-C Training Stability Observations
In a highly localized tampering scenario, CNN and ResNet model achieved high performance. However, they are sensitive to training parameter variations; their training result is not consistent, and minor variations in training configurations often resulted in significant performance differences. Therefore, they require careful hyperparameter tuning and stabilization to achieve consistent results.
VI-D Discusion
The experiment conducted demonstrate that FeatCNN-Tran, FeatCNN-TranCNN and ResNet models can effectively detect highly fragmented distortion introduced by tampering, even though such distortions are designed to mimic physiological or environmental noise. In particular, the intelligent tampering methods such 50%-50% and 75%-25% replacement, introduced subtle changes in time-domain and frequency-domain features, however this change were successfully identified by FeatCNN-TranCNN model with notable accuracy exceeding 98%.
Compared to prior studies that focus on detection of external adversarial manipulation and ECG based biometric identification, this study addresses less-explored yet critical challenge: the Intentional self-alteration and sophisticated modification of ECG signal for malicious purposes. In addition, the present work provides more advanced and robust tampering detection models by taking into account the strengths of different ML models. The study proposed improve the integrity and security of remote health monitoring systems. Table  XII summarizes the body of work reviewed in section II and the study conducted in this article. The table highlights the importance of protecting the integrity of biomedical measurements. While many of the proposed approaches use publicly available datasets generated in controlled environments to develop and test their models, some rely on self-generated datasets created outside of clinical settings. Interestingly, the performance of these approaches is not inferior to those based on datasets created in controlled environments. Our approaches fall into the second category. However, our tamper emulation and detection models are significantly more robust. Furthermore, the data acquisition process involves more complex activities, which significantly impact the quality of the ECG measurements provided to the models. Nevertheless, the accuracy they achieved for tampering detection and person identification is better than the state of the art.
TABLE XII: Compact comparison of ECG based biometric verification and tampering detection related studies

### v2

# Detection of Tampering in Wireless Electrocardiogram Using Hybrid Machine Learning

@@ Abstract
With the proliferation of wireless electrocardiogram (ECG) systems for health monitoring and authentication, protecting signal integrity against tampering is becoming increasingly important. This paper analyzes the performance of CNN, ResNet, and hybrid (Transformer+CNN) models for tamper detection. Six tampering strategies, including structured segment substitutions and random insertions, are emulated to simulate real-world attacks. The one-dimensional ECG signals are converted into a two-dimensional representation in the time-frequency domain using the continuous wavelet transform (CWT). The models are trained and evaluated using ECG data from 54 individuals recorded in four separate batches outside of clinical settings while the individuals performed seven different daily activities. Experimental results show that the hybrid models achieved an accuracy of over 99.5% in highly fragmented tampering scenarios. Even with subtle manipulations (e.g., 50%–50% and 75%–25% substitutions), the hybrid models demonstrated consistently reliable performance, achieving an average accuracy of 98%. However, the hybrid models were relatively computationally expensive.
@@ I Introduction
The electrocardiogram (ECG) plays an important role in the diagnosis and monitoring of cardiovascular diseases [1, 2]. Traditionally, measurements are performed in controlled environments by medically trained personnel. This implicitly guarantees their quality and authenticity. With growing interest in preventive medicine and long-term monitoring of cardiac patients, wireless ECGs are increasingly being used outside of clinical settings [3, 4]. This practice is gaining ground because it enables cost-effective and scalable monitoring, but it also raises concerns about quality and authenticity [5, 6]. For the measurements to be clinically relevant, these concerns must be addressed. The goal of this article is to address the issue of ECG manipulation.
In a clinical context, the ECG measurement is a legal document [7, 8]. Decisions made based on this document are binding and can have serious consequences [9]. When electrocardiograms are used outside of clinical settings, their authenticity cannot be guaranteed and they can be compromised in various ways. One of the most obvious dangers is the classic man-in-the-middle problem [10]. However, measurements can be intentionally altered by the wearer of the device for various reasons, such as to conceal or mimic certain heart conditions, to mislead employers, family members, or the legal system, or to make illegible claims against insurance policies[11, 12]. The problem is exacerbated by the fact that heart rhythms and ECG morphology change significantly when subjects move freely or exercise, making distinguishing spurious from authentic components a formidable challenge[13, 14]. This is evident in Fig. 1, which shows snapshots of a subject’s ECG while performing various exercises on the same day.
Figure 1: Snapshots of electrocardiogram measurements taken when a female subject carried out four different activities (sitting, walking, climbing stairs, and running).
In this paper, we emulate various fine-grained (less than one second in scope) manipulation strategies and propose different machine learning models—CNN-based, ResNet-based, Transformer-based, and hybrid (Transformer+CNNs)—for their detection. The models exploit correlations between salient ECG features to capture unique (for an individual) and persistent (across different activities) features in ECG measurements during highly fluctuating cardiac workloads. We train and test these models on 54 individuals performing seven different physical activities. Nineteen of our individuals were female, the remainder male. Twenty of our individuals lived in India, the remainder in Germany. The results show that the ResNet and the hybrid CNN-Transformer model achieved high accuracy in manipulation detection. The accuracy of the CNN and ResNet models in detecting highly localized manipulations is high; however, they were sensitive to variations in hyperparameters. The hybrid Transformer-CNN-based models achieved high accuracy, as well as a high degree of stability and consistent training results. However, we will show that this was associated with high computational complexity. The contributions of this article can be summarized as follows:
@@
  • Data: Our approach is real-world data-driven. A large volume of ECG data was collected from 54 subjects carrying out 7 different activities. The data were collected in four separate batches between 2019 and 2025 and exhibit a high variance, both across subjects and across activities.
  • Hybrid model for tampering detection: Multiple hybrid machine learning frameworks are proposed for tampering detection and the models are evaluated against different fine-grained tampering strategies. The hybrid models aim to capture correlations at different abstraction levels.
  • Experiments: Repeated experiments were conducted to assess the performance and stability of the models in the presence of highly variable inputs. Overall, the hybrid models achieve a detection average accuracy exceeding 98%.
@@ II Related Work
The widespread use of wireless ECG data for healthcare monitoring and person identity verification has led to concern about the wireless ECG data’s authenticity[15, 16]. Various ECG data tampering detection mechanisms are proposed to improve the reliability of wireless ECG data, and the proposed approaches are discussed in this section.
@@
The proposed approaches demonstrate reliable detection of complex signal manipulations. However, most of these approaches account for external manipulations. In contrast, the present work does not rule out internal manipulation, i.e., the deliberate manipulation of ECG data by users to achieve various goals, such as concealing heart conditions and making insurance claims. In contrast to the data used in previous case studies, which were collected in controlled environments, the present work used data collected outside of clinical settings during various activities by subjects. Such data are difficult to evaluate and contain significant motion artifacts. We propose robust machine learning models to ensure that these challenges are successfully overcome.
@@ III Data Acquisition
2019 8 8/0 30±630$\pm$ 630 ± 6 years
2024 16 5/11 27±1327$\pm$ 1327 ± 13 years
2024 10 5/5 22±1.922$\pm$ 1.922 ± 1.9 years
2025 20 17/3 24.2±2.2624.2$\pm$ 2.2624.2 ± 2.26 years
@@ V-A3 Transformer Model
V-A4 Training and Evaluation for Tampering Detection Models
@@
V-B Computational Efficiency of Models
@@
TABLE II: Model Computational Effort and Input Size in million FLOPS
Model Input Size FLOPs
@@
VI Evaluation
For tampering identification, seven models are proposed, and the performance of each model is evaluated with the six tampering scenarios. Stratified sampling is used to split the data collected into 80% training, 10% validation, and 10% testing sets. Each experiment is repeated 25 times, and the reported results represent the average accuracy over these runs.
VI-A Tampering Detection Results
As shown in Table, III, all models performed well in the classification of heavily tampered data, especially in the scenarios with sporadic 20% or 50% tampering, where detection accuracies varied from 91% to 98% respectively. The performance differences between the individual models became more apparent in more subtle cases such as 50:50 and 75:25 segment swaps. The models combining convolutional feature extraction with transformer encoding achieved some of the best results. Interestingly, ResNet was consistently competitive in all scenarios, though the transformer-based models exhibited the smallest performance variance across the 25 independent experiments. The CNN-based models exhibited significant performance fluctuations.
To further gain insights about the performance of each model, the performance matrices were separately analyzed for each model and the average values are summarized in Tables  IV. Accordingly, the FeatCNN-TransCNN and FeatCNN-Trans models obtained the highest accuracies of 99.45% and 99.37%, respectively.
TABLE IV: The average values of accuracy, precision, recall, and F1-score for all Tampering scenario
@@ VI-A Tampering Detection Results
Tran-DeepFFN 88.87 87.84 89.88 88.85
Tran-CNNFFN 90.37 89.5 91.23 90.35
FeatCNN-Tran 99.37 99.33 99.4 99.37
FeatCNN-TranCNN 99.45 99.49 99.41 99.45
CWT-FeatCNN-Tran 97.26 96.91 97.57 97.24
CNN 99.16 98.92 99.37 99.14
ResNet 99.19 98.82 99.66 99.16
VI-B Training Stability Observations
In a highly localized tampering scenario, the CNN and ResNet models achieved high performance, nevertheless, they were sensitive to training parameter variations, exhibiting inconsistent training results. Minor variations in training configurations often led to significant performance differences. Therefore, they required careful hyperparameter tuning and stabilization to achieve consistent results.
TABLE V: Compact comparison of ECG based biometric verification and tampering detection related studies
@@ VI-B Training Stability Observations
VI-C Discussion
The experiment conducted demonstrate that FeatCNN-Tran, FeatCNN-TranCNN and ResNet models could effectively detect highly fragmented distortion introduced by tampering, even though such distortions could be easily mistaken for physiological or environmental noise. In particular, the intelligent tampering methods (50%-50% and 75%-25% replacement) introduced subtle changes in time-domain and frequency-domain features, however these modifications were successfully identified by FeatCNN-TranCNN model with notable accuracy exceeding 98%. This said, the transformer-based models incurred a considerable computation cost, as can be seen in Table II.
Compared to prior studies that focus on detection of external adversarial manipulation and ECG based biometric identification, this study addresses less-explored yet critical challenge: the Intentional self-alteration and sophisticated modification of ECG signal for malicious purposes. In addition, the present work provides more advanced and robust tampering detection models by taking into account the strengths of different ML models. The study proposed improve the integrity and security of remote health monitoring systems. Table  V summarizes the body of work reviewed in section II and the study conducted in this article. The table highlights the importance of protecting the integrity of biomedical measurements. While many of the proposed approaches use publicly available datasets generated in controlled environments to develop and test their models, some rely on self-generated datasets created outside of clinical settings. Interestingly, the performance of these approaches is not inferior to those based on datasets created in controlled environments. Our approaches fall into the second category. However, our tamper emulation and detection models are significantly more robust. Furthermore, the data acquisition process involves more complex activities, which significantly impact the quality of the ECG measurements provided to the models. Nevertheless, the accuracy they achieved for tampering detection and person identification is better than the state of the art.